OpenStudy (anonymous):

The speed of sound V near the Earth's surface can be represented by this equation, v=20sqrt t +273 , where t is the temperature in degrees Celsius. Find the temperature when V = 380 meters per second. Round to the nearest whole number.

OpenStudy (amistre64):

if we solve for "t" and input V we should be able to get it

OpenStudy (amistre64):

assuming V = v that is

OpenStudy (amistre64):

\[v=20\sqrt{t}+273\] \[v-273=20\sqrt{t}+273-273\] \[v-273=20\sqrt{t}+273-273\] \[\frac{v-273}{20}=\frac{20\sqrt{t}+273-273}{20}\] \[\left(\frac{v-273}{20}\right) ^2=\left(\frac{20\sqrt{t}+273-273}{20}\right)^2\]